# Interpreting artificial neural networks to detect genome-wide association signals for complex traits

### v1

@@ Results
For 6 different phenotype simulations with different noise scaling factors ($k$ = 1, $k$ = 2, $k$ = 3) and different numbers of causal SNPs (100 and 1000 causal SNPs), we tested the performance of logistic regression (LR), saliency map (SM), integrated gradient (IG) and permutation-based (PM) feature importance approaches based on precision and ability to detect different types of causal positions (interaction and dominant/recessive). An important note here is that we defined true positive (TP) as a detected position (i.e., PAL) which is in high correlation ($r^{2}>0.5$ in 100 SNP upstream and 100 SNP downstream region, total $\sim$2Mb) with any true causal position (since high LD, which is prevalent in proximal loci, makes it difficult to detect the exact causal SNP). The results show that LR with Bonferroni correction seems to produce the least number of false positives over all different simulations but $PAL_{AMAS}$ and $PAL_{Common}$ obtained from the IG approach with strict thresholding (99.99 percentile) was a close contender (S1 Fig). Interestingly but perhaps expectedly, the LR approach predominantly detected the causal positions with interactive genetic effects, whereas the highest percentage of dominant/recessive positions were detected by gradient-based approaches (Table 1). We furthermore visualised the performance of IG on Manhattan plots obtained from LR (S2 Fig). Even under high noise schemes, some high-effect positions were detected both by LR and our method.
Table 1: Percentage of dominant/recessive positions (as opposed to interactive positions) in correctly detected causal loci averaged over all 6 simulation scenarios with varying parameters. Significance-based detection refers to Bonferroni correction for LR (logistic regression) and strict threshold (99.99 percentile) for IG (integrated gradients), SM (saliency map) and PM (permutation-based) approaches. Since significance thresholds are not directly comparable between LR and other methods, percentages were also provided for the top 5 and top 10 correctly detected causal loci. The expected frequency for dominant/recessive positions based on the phenotype simulations is 0.5. Values in parentheses demonstrate counts as [dominant/recessive]/([dominant/recessive] + [interactive]).
Category Top 5 Top 10 Significance-based
LR 0.0000 (0/24) 0.0213 (1/47) 0.0902 (12/133)
IG 0.1304 (3/23) 0.2143 (9/42) 0.1268 (9/71)
SM 0.0000 (0/23) 0.1500 (6/40) 0.1549 (11/71)
PM 0.0500 (1/20) 0.1579 (6/38) 0.1618 (11/68)
@@
After testing our approaches on simulations, we applied the IG method on the EstBB SCZ cohort to detect PAL. Although it is difficult to compare the methods reliably (see Discussion), we decided not to use SM and PM on real data due to higher false positives under certain simulation scenarios (S1 Fig). We detected a single locus with the strict threshold and multiple loci with the relaxed threshold in genomic regions with multiple protein-coding genes (Fig 2). The highest confidence signal on chromosome 4 (PAL d) included variants within TET2 and PPA2 genes and rs10010325 SNP, which has been reported to have pleiotropic effects for attention deficit hyperactivity disorder, autism spectrum disorder and intelligence [25]. Furthermore, in this LD region resides rs7674220 (albeit not being present in our non-imputed array framework), which is associated with SCZ and mapped to TET2 [26]. Interestingly, rs2647259 present in this region in our data seems to be associated with PPA2 expression levels in brain tissues (based on variant-gene association analysis, explained in the next paragraph). Other high-confidence regions we detected in chromosomes 3 (PAL c) and 11 (PAL f) are gene-dense regions with multiple low p-value SNPs in Trubetskoy et al. metadata (i.e., largest SCZ GWAS meta-analysis to date) [27]. The high signal LD region we detect in chromosome 11 includes rs1892928 (not present in our non-imputed array data), which was reported to be associated with SCZ and mapped to SYT12. In this region, we also detected rs7122539 (present in our array framework), which was shown to be associated with bipolar disorder and mapped to PC gene [28, 29]. The region in chromosome 3 includes rs4955417 (associated with depressive symptoms and neuroticism, mapped to IHO1 and C3orf62 [30, 31]) and rs7617480 (associated with major depressive disorder and unipolar depression [32, 33, 34]). Furthermore, QRICH1 gene reported in the Schizophrenia Exome Sequencing Meta-Analysis (SCHEMA) with p-value$<$0.01 and reported to be associated with developmental disorders and autism spectrum disorder resides in this region [35, 36, 27, 37]. Other lower confidence but potentially interesting regions include C1orf87 (PAL a) (with gene ontology annotation calcium ion binding), and AOX1 (PAL b) (with gene ontology annotations oxidoreductase and electron transfer activity) genes. SNPs in detected PAL are presented in S1 Table, genes in detected PAL are presented in S2 Table, and SNPs with low p-values ($<$0.001) from Trubetskoy et al. 2022 metadata in detected high LD regions are presented in S3 Table.
@@
Figure 2: PAL detected by integrated gradients (IG) approach. Red dashed lines indicate significance thresholds (relaxed and strict) and blue markers indicate PAL above threshold over all trained 10 models (i.e., $PAL_{Common}$). For all PAL (a-g), protein coding genes in those regions were provided except for multi-gene regions ($>$2 genes) for which the most likely mapped gene based on literature was provided inside the parentheses.
We further assessed whether variants in the PAL we detected were enriched in genes with brain associations. For this purpose, we used the open access Adult Genotype Tissue Expression (GTEx) Project [38] variant-gene association dataset (GTEx_Analysis_v8_eQTL). Since we did not fine map the causal variants, we assessed the significant expression levels (significance defined as pval_nominal $<$ pval_nominal_threshold in significant variant-gene pairs dataset from GTEx *.signif_variant_gene_pairs.txt.gz files) in all brain tissues for all detected loci and close proximity loci (100 SNP upstream and 100 SNP downstream region of the PAL, total $\sim$2Mb) with high LD ($>$0.5). We initially obtained the empirical cumulative distribution function (ECDF) for significant gene-tissue (only brain tissues) expression change in all 290,522 positions. Mean value for all positions was 0.858 (i.e., on average, a SNP had 0.858 gene-tissue combinations with significantly altered expression in brain) with most positions (250,692 SNPs) having no variants significantly affecting gene expression in brain tissues. Percentile ranks for all methods detecting PAL was higher than 90, indicating potential enrichment in brain expressed genes in these detected regions (S3 Fig). As an additional step, we used the GENE2FUNC tool in FUMA [39] to assess functionality for the set of genes present in the detected PAL. In terms of expression, there was significant downregulation in brain tissues and in terms of GWAS enrichment, lowest enrichment p-values were related to brain morphology and volume with multiple involved genes (S4 Fig). It is important to note here that we directly tested all the genes in PAL without any gene mapping attempt, with most genes coming from gene-rich loci from chromosomes 3 and 11 (PAL c and f), which might potentially be introducing bias to these results. Another potential biasing caveat for both ECDF and functional analysis is that detected PAL generally consist of multiple SNPs in high LD, and it is likely that only a small portion of SNPs in these LD blocks indicate a true signal.
@@
We furthermore compared these approaches with conventional GWAS using LR and REGENIE association analysis on the real dataset. Although there was a high correlation between MAS and -log(p) values (S5 Fig), there was no significant signal detected using GWAS, and only a single PAL overlap was present in the top 10 signals (S6 Fig).
@@ Discussion
In this work, we assessed post hoc neural network interpretability methods for detecting disease-associated genomic loci and proposed a general framework for using these methods within a genotype-phenotype context. Simulations demonstrated that these approaches can provide low false positive rates with strict thresholding even under relatively low genetic signal scenarios. An important note here is that the noise factor of 3 in simulations approximately corresponds to the SNP-based heritability estimate for SCZ, which is around 0.25 [27]. However, this estimate is based on linear additive models, whereas our simulated traits were based on interactive and dominant/recessive effect positions, possibly making detection of associated loci substantially more difficult. This brings us to a fundamental challenge in method development for causal detection, which is the lack of ground truth disease models. Without knowing the exact disease models, it is not trivial to assess these methods due to the likely discrepancy between simulated and real domains. In this regard, a potential solution, knockoff variables, has been adopted for neural networks in previous works [40, 13, 10]. Briefly, these are implanted “fake” positions following the distribution of the real input features to assess the false positive rates when detecting relevant features. However, sampling knockoff variables is not straightforward because deviations from the distribution of real input features would bias the false positive estimates. Furthermore, implementing a sufficient number of knockoff variables would potentially deteriorate the realness of data and significantly increase the computational burden due to the fully connected architecture of our models. Future parameter-light architectures (such as convolutions) could be more suitable to explore this approach. Another challenge for method assessment arises due to LD. Even through simulations with ground truth information, evaluation metrics such as TP/FP counts can be substantially biased since methods (including LR) tend to present highly correlated signals. With our definition of TP (i.e., a detected position in high correlation with any true causal position), a method might present higher TP (or FP) counts simply due to multiple signals in a high LD block as opposed to another method with signals consisting of isolated SNPs. Another LD-related challenge (but specific to our application) is due to models trained with different seeds detecting different signals in high LD. To increase our power, we counted these high LD PAL as common signals between models, which boosted TP signals in simulation analyses (S1 Fig). Although a substantial portion of this boost was probably due to overlapping signals, the same TP/FP analysis where we only considered exact detected SNPs as common between models demonstrated that our initial approach (common signals based on LD) allows catching more TP signals and FP signals do not change substantially between two approaches for the IG method (S7 Fig). Considering all these issues, future research might benefit from evaluation metrics better tailored for genomic data.
@@
Another important subject related to neural networks is stochasticity. In our approach, we employed both heavy dropout regularisation (including unusual dropout masking before the input layer) and ensemble-like multiple seed training to take advantage of stochasticity and limit false positives. Interestingly, even with this heavy regularisation, we were able to train models successfully (i.e., positive correlation between training and validation loss), even though the predictive power of these models was limited. Receiver operating characteristic (ROC) curves for phenotype prediction in the real dataset were comparable between neural network and logistic regression models, with logistic regression being marginally better (S8 Fig). Initially, this might suggest that the real dataset primarily exhibits linear effects. Yet, as a counter-evidence to the “mainly linear effects” scheme, logistic regression could not detect any significant positions when applied to the real cohort, and there was no correspondence between the top detected positions of logistic regression and neural network approaches (S6 Fig), which might suggest the existence of nonlinearities in the real dataset. These aspects of effect types and how heavy regularisation alters the capacity of neural networks for detecting nonlinearities are interesting subjects for future research. Considering the relatively small dataset for a case/cohort type analysis in this work, increasing the training data and experimenting with more simulation scenarios (such as including more nonlinearities, gene-environment interactions or covariate effects) could be helpful to distinguish the differences between linear and nonlinear approaches.
@@
An intriguing finding was that all approaches mainly detected positions with interactive effects while neural network methods detected more positions with dominant/recessive effects compared to LR. This is potentially due to dominant/recessive positions being more difficult to detect since the SNP effect completely disappears for a proportion of genotypes (see Materials and Methods). Another related point is that the outcome space of functions defining the disease genomics is considerably smaller due to the ternary nature of the genotype variables. Research on the topology of this space could provide valuable insights on limitations of methods developed for in silico detection of causal loci and provide better roadmaps during development.
@@
Despite the state-of-the-art deep learning algorithms and the ever-increasing amount of data in biobanks, there are still multiple challenges to the wide-spread adoption of neural networks for GWAS, such as stochasticity stemming from initial random states, lack of established uncertainty quantification and lack of ground truth disease models. These challenges are also fundamentally connected to very interesting questions, such as the contribution of nonlinear effects (gene-gene or gene-environment interactions or other polynomial disease models) or omnigenic vs polygenic models for complex disease development [41]. In the near future, assumption-free interpretable neural networks can become important tools to explore these questions and potentially replace linear models as the conventional method of conducting GWAS and obtaining interpretable polygenic risk scores.

### v2

@@ Results
As post hoc approaches which can be applied to any trained neural network models, we evaluated the performance of the saliency map (SM), integrated gradient (IG) and permutation-based (PM) feature importance methods across six different phenotype simulations, which varied by noise scaling factors ($k$ = 1, $k$ = 2, $k$ = 3) and the number of causal SNPs (100 and 1000). We designated a signal as true positive (TP) if one or more SNPs within a detected PAL block (i.e., blocks formed by clumping detected SNPs less than $100$kb distance) were in close proximity (±$\pm 100$kb, approximately ±$\pm 20$ SNPs) to a causal position. This approach both accounts for the difficulty of pinpointing the exact causal SNP due to LD between proximal SNPs, and also mitigates bloated TP counts due to multiple signals from the same LD blocks. The results showed that the IG method excels over different simulation settings with the strict $\theta$ threshold in terms of controlling false positives (FPs, i.e., type I errors) compared to other methods, and also performs fairly well in low and moderate noise settings with the relaxed $\theta$ threshold (S3 Fig). Since IG had better control of FPs and also satisfied the assumption of half-normal null distribution for obtaining p-values (see Materials and Methods, S1 Fig), we compared IG to logistic regression (LR) under the same simulation settings for different significant p-value thresholds, in which 1.7e-07 is the stringent Bonferroni threshold obtained by dividing p-value = 0.05 by the total number of SNPs (Fig 3 and S4 Fig). We performed comparisons for different p-value thresholds since multiple testing correction is not straightforward for neural network models (see Discussion). The IG method controlled for FPs better than LR for most simulation parameters and significance thresholds, albeit generally presenting fewer TPs. We further compared the types of causal SNPs (dominant/recessive versus interactive) detected by these two approaches (S5 Fig). Interestingly, mainly positions with interactive genetic effects were detected and there was no apparent difference between the IG and LR method in this aspect. Finally, we visualized the performance of IG on Manhattan plots obtained from LR, demonstrating the capacity of both approaches for detecting multiple causal SNPs (S6 Fig).
Figure 2: Comparison of true positive (TP) and false positive (FP) counts for the integrated gradient (IG) and logistic regression (LR) methods based on various simulation scenarios (n: noise factor, c: number of causal positions, t: relaxed $\theta$ threshold) and different significant p-value thresholds. A signal was determined to be TP if one or more SNP in a detected PAL block (i.e., blocks formed by clumping detected SNPs less than $100$kb distance) was in close proximity (±$\pm 100$kb, approximately ±$\pm 20$ SNPs) with a causal position. Positive values (above 0 on the y-axis) indicate TP counts whereas negative values (below 0 on the y-axis) indicate FP counts.
@@
After testing the approaches using simulations, we applied the IG method on the EstBB SCZ cohort to detect PAL. We opted not to use SM and PM on real data due to their higher false positive rates under certain simulation scenarios (S3 Fig) and difficulty of obtaining p-values (see Materials and Methods). We detected a single locus with the strict $\theta$ threshold and multiple loci with the relaxed $\theta$ threshold in genomic regions with multiple protein-coding genes (Fig 3, Table 1). The signal with the highest confidence (p-val $\lessapprox 5\times 10^{-8}$) on chromosome 11 (PAL f) has been found to be associated with bipolar disorder (a genetically closely related trait with SCZ [34]), with the loci reported in literature including all the genes we report with positional and eQTL mapping [35]. The lead SNP (rs7944999) in this region is in LD ($r^{2}=0.33$) with SCZ sub-threshold SNP rs1892928 (p-value = 6.8e-07, within SYT12) reported in Trubetskoy et al. meta-analysis [36]. Furthermore, it is also in LD ($r^{2}=0.32$) with rs10896135 located in C11orf80, a 98 kb open reading frame, and rs7122539 ($r^{2}=0.46$) in PC, which are similarly associated with bipolar disorder [35, 37]. Detected PAL b is a gene-dense region found to be associated with mood disturbance (depressive, psychotic and manic symptoms, bipolar II and major depressive disorder), with almost complete overlap with our positionally and eQTL mapped genes [38]. This region includes rs4955417 ($r^{2}=0.42$ with PAL b lead SNP), which has been associated with depressive symptoms and neuroticism, and was mapped to IHO1 and C3orf62 [39, 40]), and rs7617480 ($r^{2}=0.33$ with PAL b lead SNP), linked with major depressive disorder and unipolar depression, and mapped to KLHDC8B [41, 42, 43]. Furthermore, QRICH1 gene within this PAL and reported in the Schizophrenia Exome Sequencing Meta-Analysis (SCHEMA) with p-value$<$0.01, has been linked to developmental disorders and autism spectrum disorder. This region also includes IMPDH2 and IP6K2, both reported in SCHEMA with p-value$<$0.05 [44, 45, 36, 46]. The PAL d region on chromosome 4 contains multiple variants linked to neuropsychiatric disorders. For instance, rs7674220 ($r^{2}=0.28$ with PAL d lead SNP) has been associated with SCZ and mapped to TET2 [47]. Additionally, rs10010325 ($r^{2}=0.48$) (mapped to TET2), rs2726528 ($r^{2}=0.48$), rs2713871 ($r^{2}=0.36$) and rs1603705 ($r^{2}=0.20$) (all mapped to PPA2), which have been reported to have pleiotropic effects for attention deficit hyperactivity disorder, autism spectrum disorder and intelligence [48]. TET2 gene was also reported in SCHEMA with p-value$<$0.05. An interesting signal we detected is PAL g on chromosome 12, which we positionally mapped to NTF3. NTF3 encodes Neurotrophin-3 protein which has been shown to be linked to SCZ in multiple studies [49, 50], but with no reported SNP in GWAS Catalog for SCZ.
In addition, we assessed the aggregate protein-coding genes we detected in terms of enrichment in phenotypes based on GWAS Catalog (S7 Fig). Lowest enrichment p-value phenotypes were brain morphology and volume, with multiple genes also linked with depressive symptoms, mood instability and intelligence. Among detected variants, rs7647812 (PRKAR2A), rs6779394 (USP19), rs7944999 (PC), and rs10461139, rs10010325, rs6855629, rs2454206 (TET2) are targets of nonsense-mediated mRNA decay, while rs4955418 (CCDC71) and rs2454206 (TET2) are missense variants. The extended information about detected PAL and SNPs is provided in S1 Table.
Table 1: Detected PAL with lead SNPs and mapped protein coding genes. Genes encompassing lead SNPs are marked with a star.
PAL Lead SNP Chr:Pos P-val Positional Mapped (±100 kb) eQTL Mapped
a rs4915842 chr1:59988397 $1.2\times 10^{-5}$ C1orf87*, CYP2J2 NaN
b rs6779394 chr3:49120338 $7.8\times 10^{-6}$ C3orf62, C3orf84, CCDC71,
DALRD3, GPX1, IHO1,
IMPDH2, IP6K2,
KLHDC8B, LAMB2,
NDUFAF3, PRKAR2A,
QARS1, QRICH1, RHOA,
SLC25A20, USP19*, USP4
AMT, CCDC71, DALRD3,
GMPPB, GPX1,
KLHDC8B, NCKIPSD,
NICN1, P4HTM, QRICH1, WDR6
c rs12642383 chr4:28192894 $3.5\times 10^{-5}$ NaN NaN
d rs10461139 chr4:105184142 $3.4\times 10^{-6}$ PPA2, TET2* NaN
e rs7454792 chr6:82072955 $1.9\times 10^{-5}$ IBTK NaN
f rs7944999 chr11:66920392 $\lessapprox 5\times 10^{-8}$ C11orf80, C11orf86,
LRFN4, PC*, RCE1, SYT12
PC
g rs11063650 chr12:5403667 $3.4\times 10^{-5}$ NTF3 NaN
@@
Figure 3: PAL detected by the integrated gradients (IG) approach. Red dashed lines indicate $\theta$ thresholds (relaxed and strict) and blue markers indicate PAL above threshold over all trained 10 models (i.e., $PAL_{Common}$). The closest protein coding gene to the lead SNP in ±$\pm 100$kb region was provided for PAL (a-g).
We further assessed whether variants in detected PAL were enriched in genes expressed in brain. We initially obtained the empirical cumulative distribution for significant gene-tissue (only brain tissues) expression change in all 290,522 positions. The mean value for all positions was 0.858 (i.e., on average, a SNP had $<1$ gene-tissue combinations with significantly altered expression in brain) with most positions (250,692 SNPs) having no variants significantly affecting gene expression in brain tissues. However, the percentile rank for the detected PAL was 99.4, indicating potential enrichment in brain expressed genes in these detected regions (S8 Fig). As an additional step, we assessed expression specificity for the set of genes within the detected PAL and found significant downregulation in multiple brain tissues (S9 Fig).
@@
Finally, we compared our approach to conventional GWAS using LR and REGENIE association analyses on the real dataset. Although there was high correlation between MAS and -log(p) values (S10 Fig), there was no significant signal detected using LR or REGENIE, and no overlap in signals at lower p-value significance thresholds (S11 Fig). These conventional GWAS results are in line with findings from the UK Biobank, which used a similarly sized case sample set and reported no genome-wide significant hits [51].
@@ Discussion
Method assessment and simulations. In this work, we assessed post hoc neural network interpretability methods for detecting trait-associated genomic loci and proposed a general framework for using these methods within a genotype-phenotype context. Simulations demonstrated that these approaches can yield low FP rates with a strict $\theta$ threshold, even in scenarios with relatively weak genetic signals. In addition, the method we proposed for obtaining p-values was robust not only to different scenarios but also to different $\theta$ thresholds for controlling FPs (Fig 3 and S4 Fig). An important point to note is that the noise factor of 3 in the simulations corresponds approximately to the SNP-based heritability estimate for SCZ, which is around 0.25 [36]. However, this estimate is based on linear additive models, whereas our simulated traits were based solely on interactive and dominant/recessive effect positions, possibly making detection of associated loci substantially more difficult. Indeed, the neural network model was most successfully trained with the real dataset (i.e., smallest validation loss and least overfitting on average), which in addition increases confidence for the findings from the real dataset (S12 Fig). This connects to a fundamental challenge in method development for detecting causal variants - the absence of ground truth of the underlying model for a trait (i.e., genetic architecture and the interactive landscape of genetic and environmental effects). Without knowing the underlying model, assessing these methods is challenging due to the likely discrepancy between simulated and real data domains. Another challenge for method assessment arises from LD. Even in simulations with ground truth information, evaluation metrics such as TP/FP counts can be substantially biased since methods (including LR) tend to produce highly correlated signals. We aimed to mitigate this with our definition of TP (i.e., if one or more SNP in a detected block is in ±$\pm 100$kb region of a causal position), but methods with different capacities for detecting signals in LD regions might still be difficult to compare reliably. Future research could benefit from evaluation metrics better tailored for genomic data.
Uncertainty quantification. An important obstacle to the broader adoption of the proposed approaches for GWAS lies in the difficulty of establishing reliable statistical significance measures. While our approach for obtaining p-values effectively controlled FPs in difficult simulation settings, a stable multiple testing correction remains to be a challenge due to complex dependencies stemming from the neural network architecture and gradient calculations. Furthermore, we used a pooled approach to obtain p-values based on a global null distribution for all features, essentially testing for relative extremity. This is fundamentally different from conventional GWAS where each SNP is tested against its own null distribution. The global null approach allowed us to obtain meaningful p-values (i.e., can be corrected considering thousands of SNPs) via a feasible number of sampling iterations but it also inherently limits the detection of small but significant effects. Another important limitation is that it might not be possible to estimate the null distribution for all types of feature importance scores (S1 Fig), which is especially relevant if a complete permutation approach with multiple training runs is not computationally feasible as in our application. As a conceptually different solution for controlling FPs, knockoff variables have been adopted in previous works [52, 13, 10]. Briefly, these are implanted “fake” positions which mimic the distribution of real input features and enable the estimation of FP rates due to their independence from the output variable. However, sampling knockoff variables is not straightforward because deviations from the distribution of real input features would bias the false positive estimates. Furthermore, implementing a sufficient number of knockoff variables would potentially deteriorate the realness of data and significantly increase the computational burden due to the fully connected architecture of our model. Future parameter-light architectures (such as convolutions) or prior feature selection could be more suitable to explore this approach.
@@
Linear versus non-linear models. Another important subject related to neural networks is model training. In our approach, we employed both heavy dropout regularisation (including unusual dropout masking before the input layer) and ensemble-like multiple seed training to take advantage of stochasticity and limit FPs consequently. Interestingly, even with this heavy regularisation, we were able to train models successfully (i.e., positive correlation between training and validation loss), even though the predictive power of these models was limited. Receiver operating characteristic (ROC) curves for phenotype prediction in the real dataset were comparable between neural network and logistic regression models, with logistic regression being marginally better (S13 Fig). Initially, this might suggest that the real dataset primarily exhibits linear effects. Yet, as a counter-evidence to the “mainly linear effects” scheme, logistic regression or REGENIE could not detect any significant positions when applied to the real cohort, and there was no correspondence between the top detected positions of linear and neural network approaches (S11 Fig), which might suggest the existence of nonlinearities in the real dataset. These aspects of effect types and how heavy regularisation alters the capacity of neural networks for detecting nonlinearities are interesting subjects for future research. Considering the relatively small dataset for a case/cohort type analysis in this work, increasing the training data and experimenting with more simulation scenarios (such as including more nonlinearities, gene-environment interactions or covariate effects) could be helpful to distinguish the differences between linear and nonlinear methodologies.
@@
Interactive and dominant/recessive effects. One intriguing finding was that all approaches mainly detected positions with interactive effects. This may be due to dominant/recessive positions being more difficult to detect since the SNP effect completely disappears for a proportion of genotypes (see Materials and Methods). Additionally, the outcome space of functions defining complex trait genomics is considerably smaller than that of continuous variables, such as gene expression measurements, due to the ternary nature of genotype variables. Research into the topology of this space could provide valuable insights about the limitations of methods developed for in silico detection of causal loci, and provide better roadmaps and improvement strategies.
Application to real datasets. In SCZ cohort application, we presented extensive evidence based on literature and enrichment analyses to demonstrate that the detected loci are related to neuropsychiatric disorders and brain function. Future research should focus on replicating these results for different populations, biobanks and phenotypes. Furthermore, to mitigate confounding factors, we distribution matched case and control datasets based on age, sex and BMI, and pruned the data to eliminate cryptic relatedness. Since EstBB dataset do not exhibit high population structure (e.g., LR findings did not change when PCs are used as covariates, S2 Fig), we did not include PCs as input features for the neural network model. Future models could integrate potential confounding variables to improve generalizability over different datasets and reduce data preprocessing steps.
@@
Final remarks. Despite the state-of-the-art deep learning algorithms and the ever-increasing amount of data in biobanks, several challenges remain in the wide-spread adoption of neural networks for GWAS. These include stochasticity stemming from initial random states, lack of established uncertainty quantification and the absence of ground truth for the underlying genetic and environmental landscape of complex traits. These challenges are fundamentally connected to intriguing questions, such as the contribution of nonlinear effects (gene-gene or gene-environment interactions or other polynomial disease models) or omnigenic versus polygenic models for complex disease development [53]. In the near future, assumption-free interpretable neural networks may become essential tools for exploring these questions and could ultimately replace linear models as the standard approach for GWAS and the generation of interpretable polygenic risk scores.